This is part 1-4 of the same problem, not separate problems.Required Information [The following information applies to the questions displayed below.) Part 1 of 4 1.66 points Pearl E. WRequired Information Part 2 of 4 [The following information applies to the questions displayed below.] 1.66 points Pearl E. WRequired Information [The following information applies to the questions displayed below.) Part 3 of 4 1.66 points Pearl E. WRequired Information [The following information applies to the questions displayed below.] Part 4 of 4 1.66 points Pearl E. W
Answer #1

1) Allowance for uncollectible accounts

Days Amount of uncollectible
Not yet due 23000*5% = 1150
0-90 days 14300*15% = 2145
91-180 days 9300*20% = 1860
Over 180 days 16400*85% = 13940
Total 19095

2) Adjusting entry

Date account and explanation Debit Credit
Bad debt expense (19095-3300) 15795
Allowance for doubtful accounts 15795

3) Journal entry

Date account and explanation Debit Credit
July 19 Allowance for doubtful accounts 6300
Account receivable 6300

4) Journal entry

Date account and explanation Debit Credit
Sep 30 Account receivable 6300
Allowance for doubtful accounts 6300
cash 6300
Account receivable 6300